Mathematics 241 – Multivariable Calculus Quiz 2 – September 14, 2007

Your Name:

A) (5) Find the equation of the plane in R^{3} containing (1, 1, 1) and perpendicular to the line given by

\begin{aligned}x & = 1 − 3t \\ y & = 2 + 2t \\ z & = −5t \end{aligned}

B) (5) What are the rectangular (Cartesian) coordinates of the point in R^{3} with cylindrical coordinates (r, θ, z) = (5, \frac{π}{4} , 1)?